OpenStudy (anonymous):

If a runner does 1,500 J of work over 10.0 s, then how much power does she have? 150 W , 1000 W , 1500 W , or 15000 W ? :/

OpenStudy (michele_laino):

it is simple, since the power P, is given by the subsequent formula: \[P = \frac{L}{{\Delta t}}\]

OpenStudy (michele_laino):

namely: \[P = \frac{L}{{\Delta t}} = \frac{{1500}}{{10}}\]

OpenStudy (anonymous):

so 150? that is our solution? 150 W ?

OpenStudy (michele_laino):

yes! that's right!

OpenStudy (anonymous):

yay!! :) thanks!

OpenStudy (michele_laino):

thanks!! :D